# Learning Encoding-Decoding Direction Pairs to Unveil Concepts of Influence in Deep Vision Networks

### v1

@@ 7.10 Application: Counterfactual Explanations
Figure 21: Counter-factual explanations via manipulating the patch embeddings in the representation space using directional concept arithmetic. From Left to Right: 1) original image, 2) segmentation map obtained by applying the learned concept detectors on the original image representation. The image is correctly predicted as driveway with confidence $69\%$. 3) Manipulate the latent representation and replace the road with the encoding of a building. Now the image is classified as apartment-building-outdoor with confidence $22\%$. 4) Manipulate the latent representation and replace the (whole) building with the encoding of an airplane. The manipulated image is classified as runway with confidence $57\%$.
In this section, we provide a concrete example of how the learned Encoding-Decoding Direction Pairs can be harnessed in order to provide counter-factual explanations. In this case study, we will be studying ResNet18 trained on Places365 and EDDPs learned with CFM and $I=512$. We consider the input image depicted in Fig. 21 - left. This image is correctly predicted by the network as driveway with confidence $69\%$. The second image of the same figure depicts a segmentation map of that image based on the learned concept detectors. To keep the visualization simple, we only show the most dominant detected concepts. We explore two interventions in the representation space: a) concept removal: encoding the absence of a particular concept in a patch embedding b) concept addition: encoding the presence of a particular concept in a patch embedding. For both cases we make use of (20).
To compute target projected values $t$ we use the training split of Broden. When we do concept removal, we choose $t$ to be equal to the lowest $0.005\%$ quantile of projected feature activations on filter ${\bm{w}}_{i}$. Similarly, when we do concept addition, we choose $t$ to be equal to the top $0.005\%$ quantile of projected feature activations. This process ensures that the signal value of ${\bm{x}}_{{\bm{p}}}$ for concept $i$ after the intervention matches the signal value of samples with or without the concept.
We consider two counter-factual scenarios. In the first, we study how the prediction would change if the road was replaced by a building, and second, how the same prediction would change if the building was replaced by an airplane. To accomplish that, for each scenario, we perform two interventions one after the other: a concept removal followed by a concept addition. In the first scenario, the prediction is changed to apartment-building-outdoor with confidence $22\%$, while in the second, the prediction is changed to runway with confidence $57\%$. (Fig. 21). In both cases, the change in the prediction outcome aligns with human intuition.
@@ 7.11 Application: Toy Model Correction
In this experiment we demonstrate how the proposed approach may be utilized to correct a toy model that relies on controlled confounding factors to make its predictions. For the purposes of this toy experiment we use a small convolutional neural network with 5 Conv2d layers each one followed by a ReLU activation. The top of the network is comprised of a Global Average Pooling (GAP) layer and a linear head. After each convolutional layer, except the last, there is a Dropout layer with $p=0.3$. All Conv2d layers have kernel size 3x3 and stride 2 except the last one which has stride 1. Furthermore, the latent space dimensionality is set to 16 for all convolutional units. We consider the task of predicting the chess piece name from an image depicting the piece. We use the Chess Pieces dataset from Kaggle  [^2] Chess Pieces Dataset (85x85): https://www.kaggle.com/datasets/s4lman/chess-pieces-dataset-85x85 which contains a collection of images depicting chess pieces from various online platforms (i.e., piece images appearing in online play). The spatial resolution of those images is 85x85. For simplicity, we consider 3 chess pieces to be classified by the network, namely: bishop, knight, rook, thus the network predicts $K=3$ output classes. The total number of images in the dataset is $210$, 67 for bishop, 71 for knight and 72 for rook. We make a stratified train-test split with the training set ratio set to 0.7. To encourage the model to learn a bias to make its predictions, we poison half of the rook images of the training set with the watermark text “rook” on the top left of each rook image. With the introduction of this bias on half of the images, we expect that the network learns that the watermark concept has a positive influence on the rook class, while not being the only feature of positive evidence for the same class, since we include rook images in the training set without the watermark.
7.11.1 Network Training and Evaluation
We train the network with cross-entropy loss and the Adam optimizer with learning rate $0.005$ for $1000$ epochs. In the (poisoned) training set, the model achieves 100% accuracy. For evaluation, we construct three datasets based on the test split that we created earlier. First, we consider a clean test set, a dataset comprised of test images without any watermarks (Clean). Second, we consider the previous clean set but with all the images being poisoned with the watermark (Poisoned) and c), we consider the union of the previous two datasets (Clean & Poisoned). Table 12 summarizes the performance of the network in each one of the three datasets. As evidenced by the Poisoned section of the detailed confusion matrix, the watermark is a strong feature that, whenever it is present in the image, directs the prediction towards the rook class.
@@ 7.11.2 Direction Learning for Watermark Identification
We now consider the application of the proposed method in identifying the watermark direction, from the bottom up, without relying on annotations. As we’ve shown in the previous sections, the proposed approach is unsupervised and is able to identify directions influential to the model’s predictions. Since we already identified that the watermark actually influences the predictions of the network in a consistent manner, we seek to answer the following two research questions: a) can the proposed EDDP method identify the watermark as a concept? That is, is any of the learned classifiers responsible for detecting the watermark? b) Supposing the answer to (a) is positive, given the watermark’s concept detector and the respective learned signal vector, can we fix the network in order to not rely on the watermark for its predictions ?
Direction Learning We consider the last convolutional layer as our layer of study. This layer has spatial dimensionality 2x2. To learn the latent directions, we apply our method by following the learning process described in section A.3 and we use the network’s training set as our concept dataset. Furthermore, for stable learning, we learn the directions with the Augmented Lagrangian loss scheme of Section 4.5. We optimize $\lambda^{fs}\mathcal{L}^{fs}+\lambda^{cur}\mathcal{L}^{cur}$ with the target constraints $\tau^{ma}=0.8,\tau^{ic}=0.01,\tau^{eac}=0.01,\tau^{mm}=8.0,\tau^{fso}=0.1$, and weights $\lambda^{fs}=2.6$ and $\lambda_{cur}=0.25$ . The most important hyper-parameter to tune is the dimensionality of the concept space $I$.
Watermark Direction Identification We found that, when learning with $I=6$, the proposed approach is able to identify the watermark direction. By using the learned classifiers as concept detectors, we are able to group the spatial features of the 2x2 image representations into clusters of the same concept. When applied on an image representation, each learned classifier produces a form of a binary label-map, with each element of the label-map indicating whether the part of the image behind the spatial representation belongs to the concept. In Fig. 22 we provide example image segmentations based on those label-maps. From the qualitative visualizations, we see that classifier 5 identifies the watermark. Although annotations were not required to learn the direction, since this is a controlled experiment and we know in which images we injected the watermark, we are able to quantify how well this classifier can detect the concept by evaluating its IoU performance on the concept dataset. We found that this classifier detects the watermark concept with IoU 0.85.
@@ 7.11.3 Influence Testing with RCAV
We use RCAV to measure the sensitivity of the model with respect to the watermark signal vector. The sensitivity scores reported by RCAV are 1-1 for the classes bishop and knight while it is $1$ for the class rook. This implies that when the image has the watermark it becomes more rook and less bishop or knight. This quantitative score aligns with our intuition regarding the watermark.
@@ 7.11.4 Signal Vector Faithfulness to the Watermark Concept
Leveraging our knowledge of which images contain the watermark, we can determine the watermark’s encoding direction by calculating the watermark’s Pattern-CAV (Pahde et al., 2024). Subsequently, we quantify the directional alignment between the Pattern-CAV, which was learned with supervision, and the signal vector, which was learned without. We found that the cosine similarity between these two is 0.99, indicating estimations that agree.
@@ 7.11.5 Model Correction by Using the Watermark’s Encoding and Decoding Directions
Let $\bm{w}$} and $b$ denote the learned parameters of the watermark concept detector and $\hat{{\bm{s}}}$ denote the respective learned signal vector. Without re-training or fine-tuning the network, we are going to suppress the watermark artifact component from the representation whenever it is detected by the concept detector. We propose the following feature manipulation strategy that we apply at the features of the last convolutional layer.
@@ 7.11.6 Evaluation of the Corrected Model
We evaluate the corrected model according to the same protocol that we did in Section 7.11.1. The results are depicted in Table 13. Compared to the performance of the original network (Table 12), we see that the corrected model: a) has comparable accuracy as the original model on the clean test set b) is significantly more accurate on images of the poisoned dataset with an absolute improvement of +31% and c) performs substantially better on the union of clean and poisoned datasets with an absolute improvement of +14%. We also compare our correction strategy to using a random manipulation direction with the same $k$ as before. (i.e. using a random vector in place of the learned signal vector in (33)). In a series of 10 trial evaluations on the Poisoned Test set, we verified that, in 9 out of 10 cases, no improvement was achieved: the classification accuracy was the same as the original model and the confusion matrix was still the same as in Table 12-Middle. The same observation holds when considering the learned filter direction in place of the learned signal vector. The remaining case demonstrated an improvement of +15% which is still inferior to +31% when using the signal vector.
@@ 8 Limitations
Our method builds on the linear representation hypothesis, assuming that concepts are encoded as directions in the latent space. While this view aligns with several prior works, alternative formulations such as multidimensional concept discovery (Vielhaben et al., 2023) or linear subspaces (Chormai et al., 2024) exist but fall outside the scope of our study. Additionally, we constrain our experiments to settings where the number of discovered concepts is smaller than the embedding dimensionality. As such, we do not compare with methods like Sparse Autoencoders (SAEs) (Lim et al., 2024; Cunningham et al., 2024; Sharkey et al., 2022) that assume an overcomplete representation. Even though we restricted our experiments as such, future studies may consider this analysis, since our method is not technically limited in this regard. While we evaluated our approach across multiple CNN architectures, we did not investigate applications to Vision Transformers (ViTs) (Caron et al., 2021; He et al., 2022), despite existing literature suggesting that similar linear assumptions may apply in that context (Kaltampanidis et al., 2025)—this remains a promising direction for future work. Moreover, unlike some unsupervised baselines that operate on class-specific image subsets constructed from model predictions, we applied our method to unlabeled data without class filtering, due to the lack of dense annotations in the training sets. Nevertheless, our framework could, in principle, be adapted for class-conditional analysis. Finally, our study aimed to minimize hyperparameter tuning by relying on prior work (Doumanoglou et al., 2023; 2024) and empirical preliminary experiments; however, further tuning or exploration of additional hyperparameters may yield improvements in clustering quality, interpretability, or influence (Section A.8).
@@ Appendix A Appendix
  • A.2: Theoretical Comparison with Concept-Basis Extraction
  • A.3: Direction Learning Process
  • A.4: Details for the Experiment on Synthetic Data
  • A.5: Extracting Signal Values with the Filters of Concept Detectors
  • A.6: Details on the Faithfulness Assessment of Encoding-Decoding Direction Pairs
  • A.7: Details for the Experiments on Deep Image Classifiers
  • A.8: Hyper-Parameter Study with Respect to Target Separation Margin
  • A.9: Additional Experiments on the Faithfulness Assessment of the Encoding-Decoding Direction Pairs
  • A.10: Ablation Study with respect to Filter-Signal Orthogonality Loss
  • A.11: Comparison with Unsupervised Interpretable Basis Extraction and Concept-Basis Extraction in Practical Experiments
  • A.12: More Qualitative Segmentations and Statistics for Evaluating the Interpretability of the Concept Detectors
  • A.13 More Global Model Explanations via Concept Sensitivity Testing
  • A.14 More Local Explanations with Concept Contribution Maps
  • A.15 Relation to Sparse AutoEncoders (SAEs)
@@ A.1 Unsupervised Interpretable Basis Extraction and Concept-Basis Extraction Losses
This Section complements Section 3.3 by providing further details on the loss terms of Unsupervised Interpretable Basis Extraction (UIBE) (Doumanoglou et al., 2023) and Concept Basis Extraction (CBE) (Doumanoglou et al., 2024).
Sparsity Loss ($\mathcal{L}^{s}$) (Doumanoglou et al., 2023) Based on the observation that the number of semantic labels that may be attributed to an image patch is only a fraction of the set of possible semantic labels, this loss enforces sparsity on the number of concepts that may be attributed to a patch embedding. In particular, the sparsity loss for pixel $\bm{p}$} is based on minimizing entropy and is defined as:
@@
Maximum Activation Loss ($\mathcal{L}^{ma}$) (Doumanoglou et al., 2023)
@@
Inactive Classifier Loss ($\mathcal{L}^{ic}$) (Doumanoglou et al., 2024)
@@
Maximum Margin Loss ($\mathcal{L}^{mm}$) In the original formulation of Doumanoglou et al. (2023), the Maximum Margin Loss was defined as $\mathcal{L}^{mm}=\frac{1}{M}$ with $M$ being a single parameter for the whole set of classifiers since the optimization was performed in the standardized space with shared parameters for the margins $M$ and biases $b$. In this work, we removed the standardized space constraints and instead, we have a margin parameter $M_{i}$ for each classifier in the set. Thus, we modify the Maximum Margin loss to become:
@@
A.2 Theoretical Comparison with Concept-Basis Extraction
To improve the interpretability of the discovered directions, Doumanoglou et al. (2024) made a first attempt to exploit the knowledge encoded in the network, through feature manipulation. In particular, it was suggested that uncertain network predictions shall occur when there is no positive concept content in the features. This previous approach is similar to ours in the sense that it uses feature manipulation and subsequent entropy maximization on the network’s prediction outcomes. Yet, the fundamental difference with what we propose here lies within the feature manipulation strategy. This previous approach suggested that representations ${\bm{x}}_{{\bm{p}}}$ should be manipulated towards the concept detectors’ hyperplanes, only for the concepts that are present in ${\bm{x}}_{{\bm{p}}}$ (i.e. manipulating towards the negative direction of the filter weights, when the respective concept detectors classify ${\bm{x}}_{{\bm{p}}}$ positively). Features were not manipulated in the direction of the weights to bring them towards the separating hyperplane, when they lie in the subspace of negative concept classification. While this previous attempt tried to exploit the uncertainty region of the model, it was essentially suggested that the network’s predictions should be uncertain when for all ${\bm{x}}_{{\bm{p}}}$, none of the classifiers makes positive predictions, without minding about confident negative ones. This is fundamentally different from what we propose in the present work, as here we suggest that uncertain network predictions are linked to uncertain concept information. We enforce the latter by manipulating all all features towards the hyperplanes, regardless of whether features were positively or negatively classified for the presence of a concept. We once again emphasize that here we suggest that the network’s predictions should be maximally uncertain when for all ${\bm{x}}_{{\bm{p}}}$ none of the classifiers makes confident predictions, either positive or negative. Additionally, in Doumanoglou et al. (2024), signal directions were completely overlooked.
A.3 Direction Learning Process

### v2

@@ 7.10 Application: Counterfactual Explanations
Figure 21: Counter-factual explanations via manipulating the patch embeddings in the representation space using directional concept arithmetic. From Left to Right: 1) original image, correctly predicted as driveway with confidence $69\%$ 2) segmentation map obtained by applying the learned concept detectors on the original image representation. 3) Manipulating the latent representation and replace the road with the encoding of a building. Now the image is classified as apartment-building-outdoor with confidence $22\%$. 4) Manipulating the latent representation and replace the (whole) building with the encoding of an airplane. The manipulated image is classified as runway with confidence $57\%$.
In this Section, we provide a concrete example of how the learned direction pairs can be harnessed to provide counter-factual explanations. In this case study, we will be studying ResNet18 and direction pairs learned with CFM and $I=512$. We consider the input image depicted in Fig. 21 - left. This image is correctly classified by the network as driveway with confidence $69\%$. The second image of the same figure depicts a segmentation map based on the learned concept detectors. To keep the visualization simple, we only show the most dominant detected concepts.
We consider two counter-factual scenarios. We study how the prediction would change if a) the road was replaced by a building, and b) if the building was replaced by an airplane. For each scenario, we perform two interventions in the representation space, one after the other: a concept removal (encoding concept absence in a patch embedding) followed by a concept addition (encoding concept presence in a patch embedding). In both scenarios, the intermediate concept removal does not alter the predicted class. However, in the first one, confidence is decreased to $30.5\%$, while in the second it is slightly increased to $70.4\%$. With concept addition, in the first scenario the prediction changes to apartment-building-outdoor with confidence $22\%$, and in the second, the prediction changes to runway with confidence $57\%$. (Fig. 21).In both cases, the change in the prediction outcome aligns with human intuition. More details for the intervention are provided in Section A.19.
@@ 7.11 Application: Toy Model Correction
In this experiment, we demonstrate how the proposed approach may be utilized to correct a toy model that relies on controlled confounding factors to make its predictions. For the purposes of this toy experiment, we use a small convolutional neural network with 5 Conv2d layers, each followed by a ReLU activation. The top of the network is comprised of a Global Average Pooling (GAP) layer and a linear head. We consider the task of predicting the chess piece name from an image depicting the piece. We use the Chess Pieces dataset from Kaggle  [^1] Chess Pieces Dataset (85x85): https://www.kaggle.com/datasets/s4lman/chess-pieces-dataset-85x85 , which contains a collection of images depicting chess pieces appearing in online play. For simplicity, we consider 3 chess pieces to be classified by the network, namely: bishop, knight, rook, thus, the network predicts $K=3$ output classes. To encourage the model to learn a bias in making its predictions, we poison half of the rook images in the training set with the watermark text “rook” in the top left corner of each rook image. With the introduction of this bias, we expect that the network learns that the watermark concept has a positive influence on the rook class, while not being the only feature of positive evidence for the same class, since we include rook images in the training set without the watermark. For further details regarding this experiment, refer to Section A.20.
7.11.1 Evaluation
For evaluation, we construct three datasets. First, we consider a Clean test set, a dataset comprised of images without any watermarks. Second, we consider a Poisoned test set comprised of all the images in the clean set but poisoned with the watermark, and c), we consider the union of the previous two datasets (Clean & Poisoned). Table 13 summarizes the performance of the network in these datasets. As evidenced by the Poisoned section of the detailed confusion matrix, the watermark is a strong feature that, whenever it is present, directs the prediction towards the class rook.
@@ 7.11.2 Direction Learning for Watermark Identification
Since we already identified that the watermark actually influences the predictions of the network in a consistent manner, we seek to answer the following two research questions: a) can the proposed EDDP method identify the watermark as a concept? That is, is any of the learned classifiers responsible for detecting the watermark?, b) Supposing the answer to (a) is positive, given the watermark’s concept detector and the respective learned signal vector, can we fix the network to not rely on the watermark for its predictions ?
We use the network’s training set as our concept dataset. We consider the last convolutional layer as our layer of study, which has a spatial dimensionality of 2x2. We found that when learning with $I=6$, the proposed approach is able to identify the watermark direction. In Fig. 22 we provide example image segmentations obtained via the learned concept detectors. From the qualitative visualizations, we see that classifier 5 identifies the watermark. We found that this classifier detects the watermark concept with IoU $0.85$.
@@ 7.11.3 Influence Testing with RCAV
We use RCAV to measure the sensitivity of the model with respect to the watermark signal vector. The sensitivity scores reported by RCAV are 1-1 for the classes bishop and knight , and $1$ for the class rook. This implies that, when the image has the watermark, it becomes more rook and less bishop or knight. This quantitative score aligns with our intuition regarding the watermark.
@@ 7.11.4 Signal Vector Faithfulness to the Watermark Concept
Leveraging our knowledge of which images contain the watermark, we can determine the watermark’s encoding direction by calculating the watermark’s Pattern-CAV (PCAV). Subsequently, we quantify the directional alignment between the Pattern-CAV, which was learned with supervision, and the signal vector, which was learned without. We found that the cosine similarity between these two is $0.99$, indicating estimations that agree.
@@ 7.11.5 Model Correction by Using the Watermark’s Encoding and Decoding Directions
Let $\bm{w}$} and $b$ denote the learned parameters of the watermark concept detector and $\hat{{\bm{s}}}$ denote the respective learned signal vector. Without re-training or fine-tuning the network, we are going to suppress the watermark artifact component from the representation whenever it is detected by the concept detector. We propose the following feature manipulation strategy, which we apply to the features of the last convolutional layer:
@@ 7.11.6 Evaluation of the Corrected Model
We evaluate the corrected model according to the protocol that was used in Section 7.11.1. The results are depicted in Table 14. Compared to the performance of the original network (Table 13), we see that the corrected model: a) has comparable accuracy to the original model on the clean test set, b) is significantly more accurate on images of the poisoned dataset with an absolute improvement of +31%, and c) performs substantially better on the union of clean and poisoned datasets with an absolute improvement of +14%.
We also compare our correction strategy to using a random manipulation direction with the same $k$ as before (i.e., using a random vector in place of the learned signal vector in (18)). In a series of 10 trial evaluations on the Poisoned Test set, we verified that in 9 out of 10 cases, no improvement was achieved: the classification accuracy was the same as that of the original model, and the confusion matrix remained the same as in Table 13-Middle. The same observation holds when considering the learned filter direction in place of the learned signal vector. The remaining case demonstrated an improvement of +15%, which is still inferior to +31% when using the signal vector.
@@ 8 Limitations
Our method builds on the linear representation hypothesis, assuming that concepts are encoded as directions in the latent space. While this view aligns with several prior works, alternative formulations such as multidimensional concept discovery (mcd) or linear subspaces (RelevantSubspaces) exist but fall outside the scope of our study. Additionally, we constrain our experiments to settings where the number of discovered concepts is smaller than the embedding dimensionality. As such, we do not compare with methods like Sparse Autoencoders (SAEs) (SAE1; SAE2; SAE3) that assume an overcomplete representation. Even though we restricted our experiments as such, future studies may consider this analysis, since our method is not technically limited in this regard. While we evaluated our approach across multiple CNN architectures, we did not investigate applications to Vision Transformers (ViTs) (dino; mae), despite existing literature suggesting that similar linear assumptions may apply in that context (xai-directions)—this remains a promising direction for future work. Moreover, unlike some methods from unsupervised state-of-the-art that operate on class-specific image subsets constructed from model predictions, we applied our method to unlabeled data without class filtering, due to the lack of dense annotations in the training sets. Nevertheless, our framework could, in principle, be adapted for class-conditional analysis. Finally, our study aimed to minimize hyperparameter tuning by relying on prior work (UIBE; CBE) and empirical preliminary experiments; however, further tuning or exploration of additional hyperparameters may yield improvements in clustering quality, interpretability, or influence (Section A.13).
@@ Appendix A Appendix
  • A.2: Encoding-Decoding Direction Pairs Algorithm
  • A.3: Theoretical Comparison with Concept-Basis Extraction
  • A.4: Details on Uncertainty Region Alignment
  • A.5: Details on Reading Concept Information and Intervening on their Encoding
  • A.6: Details on Using Encoding-Decoding Direction Pairs and the Regions of Uncertainty to Provide Concept-Based Local Explanations and Detailed Spatially-Aware Concept Contribution Maps
  • A.7: Direction Learning Process
  • A.8: Details for the Experiment on Synthetic Data
  • A.9: Extracting Signal Values with the Filters of Concept Detectors
  • A.10: Details on the Faithfulness Assessment of Encoding-Decoding Direction Pairs
  • A.11: Additional Experiments on the Faithfulness Assessment of the Encoding-Decoding Direction Pairs
  • A.12: Details for the Experiments on Deep Image Classifiers
  • A.13: Hyper-Parameter Study with Respect to Target Separation Margin
  • A.14: Ablation Study with respect to Filter-Signal Orthogonality Loss
  • A.15: Comparison with Unsupervised Interpretable Basis Extraction and Concept-Basis Extraction in Practical Experiments
  • A.16: More Qualitative Segmentations and Statistics for Evaluating the Interpretability of the Concept Detectors
  • A.17 More Global Model Explanations via Concept Sensitivity Testing
  • A.18 More Local Explanations with Concept Contribution Maps
  • A.19 Details on Counterfactual Explanation Example
  • A.20 Details on Toy Model Correction
  • A.21 Details on Computational Resources
  • A.22 Relation to Sparse AutoEncoders (SAEs)
@@ A.1 Unsupervised Interpretable Basis Extraction and Concept-Basis Extraction Losses
This Section complements Section 3.3 by providing further details on the loss terms of Unsupervised Interpretable Basis Extraction (UIBE) (UIBE) and Concept Basis Extraction (CBE) (CBE).
Sparsity Loss ($\mathcal{L}^{s}$) (UIBE) Based on the observation that the number of semantic labels that may be attributed to an image patch is only a fraction of the set of possible semantic labels, this loss enforces sparsity on the number of concepts that may be attributed to a patch embedding. In particular, the sparsity loss for pixel $\bm{p}$} is based on minimizing entropy and is defined as:
@@
Maximum Activation Loss ($\mathcal{L}^{ma}$) (UIBE)
@@
Inactive Classifier Loss ($\mathcal{L}^{ic}$) (CBE)
@@
Maximum Margin Loss ($\mathcal{L}^{mm}$) In the original formulation of UIBE, the Maximum Margin Loss was defined as $\mathcal{L}^{mm}=\frac{1}{M}$ with $M$ being a single parameter for the whole set of classifiers since the optimization was performed in the standardized space with shared parameters for the margins $M$ and biases $b$. In this work, we removed the standardized space constraints and instead, we have a margin parameter $M_{i}$ for each classifier in the set. Thus, we modify the Maximum Margin loss to become:
@@
A.2 Encoding-Decoding Direction Pairs Algorithm
Algorithm 1 provides, high-level, PyTorch-like pseudocode for learning the proposed Encoding-Decoding Direction Pairs.
Algorithm 1 PyTorch-Like Encoding-Decoding Direction Pairs Pseudocode
1:Input: ${\bm{X}}\in{\mathbb{R}}^{B\times D\times H\times W}$: Layer feature activations with $B$ denoting batch size.
2:    $f^{+}$: Upper-part of the studied deep network.
3:Learnable Parameters:
4:    ${\bm{W}}\in{\mathbb{R}}^{D\times I}$: Decoding directions (filters of concept detectors)
5:    $\hat{{\bm{S}}}\in{\mathbb{R}}^{D\times I}$: Encoding directions
6:    ${\bm{b}}\in{\mathbb{R}}^{I}$: Concept detector biases
7:Output: Loss to minimize
8:function learning_step($\bm{X}$})
9:  z = torch.conv2d($\bm{X}$$\bm{W}$}.T.unsqueeze(-1).unsqueeze(-1))
10:  p = z - $\bm{b}$}.reshape(1,-1,1,1)
11:  $\hat{{\bm{S}}}$ = signal_vectors($\bm{X}$},z,p) $\triangleright$ Use running stats for variance and covariance and p>0 to subsample
12:  y = torch.sigmoid(p)
13:  q = y / torch.sum(y,dim=1,keepdim=True)
14:  loss_fs = $\mathcal{L}^{fs}(\text{{q}})$
15:  loss_ma = $\mathcal{L}^{ma}(\text{{y,q}})$
[…]
17:  loss_ic = $\mathcal{L}^{ic}(\text{{y}})$
18:  loss_eac = $\mathcal{L}^{eac}(\text{{y}})$
[…]
20:  if with UFM then
21:    $\bm{X}$$^{\prime}$ = $g({\bm{X}},{\bm{W}})$
22:  else if with CFM then
[…]
24:  end if
25:  loss_ur = $\mathcal{L}^{ur}({\bm{X}}^{\prime},f^{+})$
26:  loss_lag = Lagrangian(loss_fs,loss_ur,loss_ma,loss_ic,loss_mm,loss_eac,loss_fso)
27:  return loss_lag
28:end function
A.3 Theoretical Comparison with Concept-Basis Extraction
Concept-Basis Extraction (CBE) CBE made an initial attempt to exploit the network’s instructions to improve the interpretability of the discovered directions. This approach, like ours, used feature manipulation, exploited the network’s instructions, and maximized the entropy of the network’s prediction outcomes for the manipulated features.
However, the key difference between the approaches lies in the manipulation strategy: the prior method suggested manipulating feature representations (${\bm{x}}_{{\bm{p}}}$) toward the concept detectors’ hyperplanes only for concepts currently present in ${\bm{x}}_{{\bm{p}}}$ (i.e., moving negatively along the filter weights when the concept was positively classified). It explicitly neglected manipulation in the opposite direction for negatively classified concepts. This essentially implied that network predictions should be uncertain only when none of the concept detectors made a positive prediction, ignoring confident negative ones.
Our approach fundamentally differs: we suggest that uncertain network predictions are tied to maximally ambiguous concept information, enforcing this by manipulating all features toward the hyperplanes, regardless of whether the concept was initially classified as present or absent. This ensures the network is maximally uncertain only when none of the detectors makes a confident prediction, whether positive or negative. Furthermore, CBE completely overlooked the critical role of signal directions.
A.4 Details on Uncertainty Region Alignment
Unconstrained Feature Manipulation (UFM): In this type of manipulation ${\bm{x}}_{{\bm{p}}}^{\prime}=g_{{\bm{p}}}({\bm{x}}_{{\bm{p}}};{\bm{W}})={\bm{x}}_{{\bm{p}}}-\bm{dx_{{\bm{p}}}}$ and $\bm{dx_{{\bm{p}}}}$ such that:
with $\bm{A}$$^{+}$ denoting the pseudo-inverse of $\bm{A}$}.
Constrained Feature Manipulation (CFM): In this type of manipulation ${\bm{x}}_{{\bm{p}}}^{\prime}=h_{{\bm{p}}}({\bm{x}}_{{\bm{p}}};{\bm{W}},\hat{{\bm{S}}})={\bm{x}}_{{\bm{p}}}-\bm{dx_{{\bm{p}}}}$, with $\bm{dx_{{\bm{p}}}}$ given by:
A.5 Details on Reading Concept Information and Intervening on their Encoding
Reading Concept Information from the Embeddings Reading concept information from a patch embedding ends up in estimating the concept’s signal value that is encoded in the representation. Supposing that the related conditions are met, in Section 3.2 it was shown that the filter direction can serve as a means to extract this value. While in the data-model of Section 3.2, the extraction of the signal’s value is exact, under the multi-concept signal-distractor data model of Section 4.1, the considered latent space bias $\bm{c}$} interferes with this value. More specifically, if we suppose ${\bm{w}}_{i}:{\bm{w}}_{i}\perp\hat{{\bm{s}}}_{j},i\neq j,{\bm{w}}_{i}^{T}\hat{{\bm{s}}}_{i}=1$ and ${\bm{w}}_{i}\perp{\bm{D}}$, the inner product between the filter and the embedding results in:
Thus, when considering our extended data model, the filter can be used to extract the signal’s value but with a constant offset, independent of ${\bm{x}}_{{\bm{p}}}$. Due to the offset being constant, the difference between two projected embeddings is equal to the difference of their signal values, e.g.:
And second, the concept’s point of uncertainty. Given a query point ${\bm{x}}_{{\bm{p}}}$, the concept’s point of uncertainty $\hat{{\bm{u}}}_{{\bm{p}}}$ is the point that lies on the concept detector’s hyperplane (thus corresponding to a point of maximum concept ambiguity) and the line defined by the query point and the direction of the signal vector. More specifically, the point of uncertainty is equal to $\hat{{\bm{u}}}_{{\bm{p}}}={\bm{x}}_{{\bm{p}}}-\kappa_{{\bm{p}},i}\hat{{\bm{s}}}_{i}$ with $\kappa_{{\bm{p}},i}\in{\mathbb{R}}$ corresponding to the solution of the following system of equations:
Since (31) corresponds to the function of the concept detector, this difference is positive whenever the concept is present in ${\bm{x}}_{{\bm{p}}}$ and negative whenever the concept is absent. Thus, centering the signal value of a patch embedding around its point of uncertainty can be accomplished by computing $\kappa_{{\bm{p}},i}$ using (30).
Intervening on Concept Encoding Altering the concept content that is encoded in a patch embedding ${\bm{x}}_{{\bm{p}}}$, ends up overwriting the signal value of the concept with a target value of interest. The target value is meant to be copied from a target embedding ${\bm{x}}_{{\bm{p}}}^{t}$. Let ${\bm{x}}_{{\bm{p}}}^{\prime}$ denote the altered embedding after the intervention. Supposing that we aim to intervene on the value of concept $i$, ${\bm{x}}_{{\bm{p}}}^{\prime}$ becomes:
${\bm{x}}_{{\bm{p}}}^{\prime}={\bm{x}}_{{\bm{p}}}+\kappa_{{\bm{p}},i}\hat{{\bm{s}}}_{i}$ (32)
A.6 Details on Using Encoding-Decoding Direction Pairs and the Regions of Uncertainty to Provide Concept-Based Local Explanations and Detailed Spatially-Aware Concept Contribution Maps
A.6.1 Terminology and Definitions
Suppose that we want to explain the prediction of an image classifier $f$ for image $\mathcal{I}$. Let $\bm{X}$} denote this image’s representation at the penultimate layer of the network (in CNNs, this is typically the last convolutional layer before GAP). Let ${l}_{c}\in{\mathbb{R}}$ denote the network’s output logit for class $c$ regarding input $\mathcal{I}$. Typically, $c$ will be chosen such that ${l}_{c}$ is the maximum logit among the output class logits, but other choices are also valid, for instance when we want to explain why the prediction deviated from a specific class. Let also ${\bm{X}}_{b}^{m}\in{\mathbb{R}}^{H\times W\times D}$ denote a baseline point in the uncertainty region of the model. We call it baseline because this point corresponds to an artificial image representation for which the prediction of the network is highly uncertain. Let ${l}_{b}^{m}\in{\mathbb{R}}$ denote the prediction logit for class $c$ that corresponds to ${\bm{X}}_{b}^{m}$, i.e. ${l}_{b}^{m}=f^{+}({\bm{X}}_{b}^{m})$. We define the following quantity:
${l}_{e}={l}_{c}-{l}_{b}^{m}$ (34)
[…]
A.6.2 Logit Difference Decomposition in terms of Concepts
Based on the aforementioned definitions, we have:
Due to the linearity of GAP, (36) can be written as:
Based on (40), the explanation logit ${l}_{e}$ is linearly expressed in terms of patches and concepts, giving detailed, spatially-aware information for concept contributions. We split (40) into four parts. The first part corresponds to contributions of concepts included in the sample (sample concept), the second to the contribution of concepts contained in an artificial baseline point of an uncertain prediction (baseline concept), the third part to a correction factor to account for imperfect direction learning convergence, and finally, a residual that we do not explain. If the correction factor for concept $i$ was zero, concept $i$ would have a positive contribution to the explanation logit whenever the difference in contributions between the sample and the baseline is positive. In the case of imperfect learning conditions, this difference must exceed the negative of the correction factor.
In the experiments, we report Concept Contribution Maps (CCMs) where each spatial element $\bm{p}$} is equal to:
visualized as a heatmap over the original image for a given concept $i$. Finally, we refer to the quantity ${\bm{w}}_{c}^{T}\hat{{\bm{s}}}_{i}$ as the Concept-Class Relation Coefficient (CCRC).
A.7 Direction Learning Process